OpenStudy (anonymous):

What is the derivative of sin(xy)? Please show your individual steps if possible.

sam (.sam.):

Do you want to find the partial derivatives?

OpenStudy (anonymous):

Yes, I think with respect to d/dx. So d/dx[sin(xy)], if that makes sense.

OpenStudy (anonymous):

here it is but its backwards

zepdrix (zepdrix):

lol it's backwards taylor! XD

zepdrix (zepdrix):

As sam asked, derivative with respect to x, \(\Large \dfrac{d}{dx}\) or partial derivative with respect to x, \(\Large \dfrac{\partial}{\partial x}\) Whichhhh? D:

OpenStudy (anonymous):

Oh, I meant derivative with respect to x (d/dx), sorry!

zepdrix (zepdrix):

Ok so we take the derivative of the outer function, \(\large \sin(stuff)\) in this case :) which gives us \(\large \cos(stuff)\) right? \[\Large \left[\sin(xy)\right]'\quad=\quad \left[\cos(xy)\right]\color{royalblue}{(xy)'}\]

zepdrix (zepdrix):

The chain rule tells us to multiply by the derivative of the "stuff" inside.

zepdrix (zepdrix):

So we still need to take the derivative of this blue portion.

zepdrix (zepdrix):

We'll apply the product rule for the blue part.\[\Large \left[\cos(xy)\right]\left[\color{royalblue}{(x)'}y+x\color{royalblue}{(y)'}\right]\]

zepdrix (zepdrix):

\[\Large (x)'\quad=\quad?\]Derivative of x, with respect to x? :)

OpenStudy (anonymous):

(x)'=1, so it'd be ycos(xy) + xy'cos(xy) after distributing (?).

zepdrix (zepdrix):

Yes, if you want to distribute :) that looks correct!

OpenStudy (anonymous):

Ok, thank you all for the help!